# On the non-uniqueness of the energy-momentum and spin currents

### v1

@@ Abstract
The macroscopic energy-momentum and spin densities of relativistic spin hydrodynamics are obtained from the ensemble average of their respective microscopic definitions (quantum operators). These microscopic definitions suffer from ambiguities, meaning that, one may obtain different forms of symmetric energy-momentum tensor and spin tensor through pseudogauge transformations (or in other words Belinfante improvement procedure). However, this ambiguity can be removed if we obtain these currents using Noether’s second theorem instead of widely used Noether’s first theorem. In this article, we use Noether’s second theorem to derive (symmetric) energy-momentum tensor and (antisymmetric) spin tensor without the need of pseudogauge transformations for free Dirac massive particles with spin one-half.
@@
This raises questions about which definition to follow and the physical interpretation of the anti-symmetric components of the EMT from the perspective of GR. The important question we are having is: can a specific decomposition (or a specific pseudogauge) of total angular momentum for a given system provide a “physically” meaningful local distribution of energy, momentum, and spin? While GR asserts that energy-momentum density is measurable through geometry, this matter is currently under intense debate, particularly when considering spin degrees of freedom [10, 11].
@@
In principle, one can have any form of super-potential but the question still remains that which one describes the data or physical. This ambiguity is longstanding in both quantum field theory and gravity. At this point, it is just a choice of specifying a pseudogauge and develop the formalism of relativistic hydrodynamics with spin.
@@
where $l$ designates several fields. Employing Noether’s second theorem symmetries we obtain [3]
@@
Similarly, following the above procedure it is easy to arrive at the following spin tensor definition
@@
$-\frac{\partial\mathcal{L}_{D}}{\partial(\partial_{\lambda}\psi)}\gamma^{\nu}\gamma^{\mu}\psi+\bar{\psi}\gamma^{\nu}\gamma^{\mu}\frac{\partial\mathcal{L}_{D}}{\partial(\partial_{\lambda}\bar{\psi})}\Bigg{]}\,,$
@@
$\mathcal{S}^{\lambda,\mu\nu}$ $=$ $\frac{i}{8}\,\bar{\psi}\left\{\gamma^{\lambda},\left(\gamma^{\mu}\gamma^{\nu}-\gamma^{\nu}\gamma^{\mu}\right)\right\}\,\psi\,.$ (31)
The forms of EMT and spin tensor has been favoured in recent studies. For instance, in a novel study Ref. [47], it is shown that the energy-density quantum fluctuations, in subsystems of hot and relativistic spin-half massive particles, for Eq. (29) are very small for small system size which is not the case for other forms of EMT (GLW or HW). For large system sizes, energy-density quantum fluctuations obtained from Eq. (29) approaches faster towards thermodynamic limit.
In a recent study, Ref. [24], it is shown that the spin tensor, Eq. (31), defines the spin operator that satisfies SO(3) angular momentum algebra. It is also shown that the spin tensors (GLW or HW) obtained through pseudogauge transformations do not fulfill angular momentum algebra SO(3). It might be possible that any definition of EMT and spin tensor obtained through the use of super-potentials do not, in general, satisfy such algebra. However, this requires further investigations to have any solid conclusion.
@@
We found that the derived EMT is symmetric in its indices and spin tensor is completely antisymmetric. Symmetric EMT makes the spin tensor conserved independently of total angular momentum conservation. Noether’s second theorem is not very widely known, however, using it, especially within the context of relativistic spin hydrodynamics, may remove the ambiguity of pseudogauge transformations and fix the super-potential. Though the EMT for the free fermions has been known, the spin tensor derivation using Noether’s second theorem is mentioned for the first time in this article. The method of derivation suggested in this paper could be useful to find suitable decompositions of momentum and spin for the comparison with other decompositions such as Ji [48] and Jaffe-Manohar [49].

### v2

@@ Abstract
The macroscopic energy-momentum and spin densities of relativistic spin hydrodynamics are obtained from the ensemble average of their respective microscopic definitions (quantum operators). These microscopic definitions suffer from ambiguities, meaning that, one may obtain different forms of symmetric energy-momentum tensor and spin tensor through pseudogauge transformations (or in other words Belinfante improvement procedure). However, this ambiguity can be removed if we obtain these currents using Noether’s second theorem instead of widely used Noether’s first theorem. In this article, we use Noether’s second theorem to derive energy-momentum and spin currents without the need of pseudogauge transformations for free Dirac massive particles with spin one-half.
@@
This raises questions about which definition to follow and the physical interpretation of the antisymmetric components of the EMT from the perspective of GR. The important question we are having is: can a specific decomposition (or a specific pseudogauge) of total angular momentum for a given system provide a “physically” meaningful local distribution of energy, momentum, and spin? While GR asserts that energy-momentum density is measurable through geometry, this matter is currently under intense debate, particularly when considering spin degrees of freedom [10, 11].
@@
In principle, one can have many forms of super-potential but the question still remains that which one describes the data or physical. This ambiguity is longstanding in both quantum field theory and gravity. At this point, it is just a choice of specifying a pseudogauge and develop the formalism of relativistic hydrodynamics with spin.
@@
where $l$ designates several fields. Employing Noether’s second theorem symmetries we obtain [4, 3]
@@
Similarly, following the above procedure we find that the spin tensor is actually vanishing
@@
$-\frac{\partial\mathcal{L}_{D}}{\partial(\partial_{\lambda}\psi)}\gamma^{\mu}\gamma^{\nu}\psi+\bar{\psi}\gamma^{\mu}\gamma^{\nu}\frac{\partial\mathcal{L}_{D}}{\partial(\partial_{\lambda}\bar{\psi})}\Bigg{]}\,,$
@@
$\mathcal{S}^{\lambda,\mu\nu}$ $=$ $0\,.$ (31)
The form of EMT, (29), has been favoured in a recent study [47] where it is shown that the energy-density quantum fluctuations, in subsystems of hot and relativistic spin-half massive particles, are very small for small system size which is not the case for other forms of EMT (GLW or HW). For large system sizes, energy-density quantum fluctuations obtained from Eq. (29) approaches faster towards thermodynamic limit.
In a recent study, Ref. [24], it is shown that the spin tensor, Eq. (31), trivially satisfies SO(3) angular momentum algebra. It is also shown that the spin tensors (GLW or HW) obtained through pseudogauge transformations do not fulfill angular momentum algebra SO(3). It might be possible that any definition of EMT and spin tensor obtained through the use of super-potentials do not, in general, satisfy such algebra. However, this requires further investigations to have any solid conclusion.
It turns out that using Noether’s second theorem we can naturally obtain EMT and spin tensor equivalent to classical GR currents.
@@
We found that the derived EMT is symmetric in its indices and spin tensor is vanishing. Noether’s second theorem is not very widely known, however, using it, especially within the context of relativistic spin hydrodynamics, may remove the ambiguity of pseudogauge transformations and fix the super-potential. Though the EMT for the free fermions has been known, the spin tensor derivation using Noether’s second theorem is mentioned for the first time in this article. The method of derivation suggested in this paper could be useful to find suitable decompositions of momentum and spin for the comparison with other decompositions such as Ji [48] and Jaffe-Manohar [49].